OpenStudy (anonymous):

how can I get the derivative of x^3 squareroot of x

OpenStudy (anonymous):

just to confirm, does it look like this?

OpenStudy (johnweldon1993):

how else can you write √x? *think of exponents

OpenStudy (anonymous):

\[x ^{3}.\sqrt{x}=x ^{3}.x^{\frac{ 1 }{ 2 }}=x^{(3+\frac{1}{2})}=x^{\frac{7}{2}}\] then, from that final form, you can just precede with derivative rule of experiential... have fun then :)